# Enhanced Computationally Efficient Long LoRA Inspired Perceiver Architectures for Auto-Regressive Language Modeling

### v1

@@ Abstract
The Transformer architecture has revolutionized the Natural Language Processing field and is the backbone of Large Language Models (LLMs). The Transformer uses the attention mechanism that computes the pair-wise similarity between its input tokens to produce latent vectors that are able to understand the semantic meaning of the input text. One of the challenges in the Transformer architecture is the quadratic complexity of the attention mechanism that prohibits the efficient processing of long sequence lengths. While many recent research works have attempted to provide a reduction from $O(n^{2})$ time complexity of attention to semi-linear complexity, it remains an unsolved problem in the sense of maintaining a high performance when such complexity is reduced. One of the important works in this respect is the Perceiver class of architectures that have demonstrated excellent performance while reducing the computation complexity. In this paper, we use the PerceiverAR that was proposed for Auto-Regressive modeling as a baseline, and provide three different architectural enhancements to it with varying computation overhead tradeoffs. Inspired by the recently proposed efficient attention computation approach of Long-LoRA, we then present an equally efficient Perceiver-based architecture (termed as Long LoRA Pereceiver - LLP) that can be used as the base architecture in LLMs instead of just a fine-tuning add-on. Our results on different benchmarks indicate impressive improvements compared to recent Transformer based models.
@@ 1 Introduction
The Transformer architecture has revolutionized the field of artificial intelligence, especially in Natural Language Processing (NLP) Vaswani (2017). The recent success of Large Language models such as ChatGPT  Achiam et al. (2023), Gemini  Team et al. (2023), Llama Touvron et al. (2023); Dubey et al. (2024), etc. with their comprehension and reasoning capabilities, is a testament to the effectiveness of the Transformer architecture. Prior to Transformers, deep Convolutional Neural Networks (CNNs) had demonstrated amazingly well results in computer vision applications, however, their performance does not show the same effectiveness when applied to NLP. One of the reasons CNNs are typically less effective in NLP is their limited receptive field, due in part to the convolution operation. The Transformer on the other hand, uses the attention mechanism. This operation measures the pairwise similarity between the words of the entire input sequence in order to comprehend it.
@@
Consider the matrix $Q$, $K$ and $V$ containing rows representing the learnt position encoded (PE) embedding of each token in $d$ dimensions (i.e., $1\times d$). Then the attention $A=\text{softmax}(QK^{T})$ contains the dot product similarity of each input token with every other token in the input sequence. For an input sequence with $n$ tokens, $Q,K\in\mathbb{R}^{n\times d}$ and attention $A\in\mathbb{R}^{n\times n}$. The Transformer divides the attention calculation into parallel heads, to be able to refine the learning and comprehend the contextual meaning of the input sequence. Each head computes the attention on a portion of the embedding dimension. The output in each head is computed by further multiplying the attention $A$ with $V$. The canonical Transformer’s operation can be summarized by the following equations Vaswani (2017), first the output of the $i^{th}$ head is:
@@
For language models, if text generation is the goal, the model is trained in an autoregressive manner where it learns to predict the next token given an input sequence of tokens. In autoregressive generation, the previously predicted token becomes part of the next input sequence. For autoregressive NLP models, the training process can be made highly effective by masking the attention matrix in a triangular fashion so that future tokens are not visible. The triangular masking helps in creating more (input, output) training pairs, as from an input training sequence of size $n$, $n-1$ training pairs can be created by simply hiding the next token one at a time. For NLP classification, the masking of the attention is not needed, as the classification decision is made on the entire input sequence.
@@
Since the attention computation in each head measures the pairwise similarity in the input sequence, its time complexity is $O(n^{2})$ if the sequence length is $n$. With larger NLP models being created operating on longer sequence lengths, and with multiple heads in each layer and many layers, the computational costs of Transformer training is becoming an important issue. Considerable research effort is being put into making the attention mechanism more efficient since it is the dominant computation. Numerous research papers have proposed ideas to reduce the quadratic time complexity of attention to linear or sub quadratic complexity. Some of the important works in this respect include: TransformerXL Dai et al. (2019), Linformer Wang et al. (2020), Longformer Beltagy et al. (2020), Reformer Kitaev et al. (2020), Performer Choromanski et al. (2020), Long-Short Attention Zhu et al. (2021), Perceiver Hawthorne et al. (2022); Jaegle et al. (2021; ), among others. Recently State Space Models Gu et al. (2021); Fu et al. ; Dao and Gu have drawn considerable research attention and provided impressive results in some domains. It is yet to be seen if they can be better alternatives for NLP generative models. Thus we limit our comparisons to transformer-based architectures in this work. We provide a brief background in the above related Transformer approaches towards achieving lower attention complexity. Then we elaborate on the PerceiverAR Hawthorne et al. (2022) architecture that we further enhance in this work.
2 Related Work and Contributions
One of the important works that reduces the quadratic time complexity of attention to $O(n)$ is Linformer Wang et al. (2020). The authors of Linformer empirically observed that attention has low rank, and therefore can be approximated by a low rank matrix. To achieve this, the $Q$ and $V$ matrices $\in\mathbb{R}^{n\times d}$ are projected to lower dimensions by using a learnable projection along the sequence dimension. This results in $Q,V\in\mathbb{R}^{k\times d}$ where $k<n$. Consequently the attention is $A=QK^{T}\in\mathbb{R}^{n\times k}$. Since $k$ is fixed the attention complexity is $O(n)$. Note that the output $A\times V$ is still $\in\mathbb{R}^{n\times d}$, i.e., input and output dimensions match. Because Linformer compresses the information along the input sequence dimension during the projection process, it cannot be used in effective autoregressive training as the masking of attention for future tokens cannot be accomplished. However, for classification problems where masking of attention is not needed, the Linformer architecture is effective in reducing attention complexity. Another notable work in the area of reducing the attention complexity is Reformer Kitaev et al. (2020) which uses locality sensitive hashing. This reduces the attention complexity to $O(n\text{log}(n))$.
A popular important architecture that allows autoregressive modeling and handles long input contexts efficiently is termed as TransformerXL Dai et al. (2019). It follows a different approach than trying to compress the input sequence. It divides the input sequence into segments and uses segment level recurrence. After processing one segment, the model reuses the hidden states from the last layer of that segment as the initial hidden state for the next segment. It also introduces relative positional encoding to allow it to handle longer contexts. TransformerXL presented excellent perplexity results on benchmarks. Its slight drawback is the recurrent mechanism that may lose information over long contexts. A mathematical approach to reducing the attention complexity is taken in Performer Choromanski et al. (2020) which uses random features and projections. The attention is decomposed into a product of non-linear functions of original query and key matrices referred to as random features. This allows the attention to be theoretically encoded more efficiently but the complexity of this encoding for long sequences may be higher.
In Longformer Beltagy et al. (2020) a different approach was employed which relies on sparse attention in order reduce its complexity. The Longformer authors proposed different sparse attention patterns such as sliding window, dilated sliding window attention and partially sparse global attention. In sliding window, only the nearby tokens are attended to, whereas in dilated sliding window, a diagonal partially sparse attention pattern is used. One architecture that partially relies on sliding window attention in handling long sequences is the long-short Transformer Zhu et al. (2021). Here the short attention refers to the sliding window while the long attention divides the entire context into small compressed segments. Both short and long attention are combined in the final attention. Since the long attention is based on compression, it may lose some important contextual information in an autoregressive generation.
Another approach to efficient attention was proposed in PerceiverAR Hawthorne et al. (2022). It accomplishes efficient handling of long contexts by diving the input sequence in two components of history and latent. The query matrix is computed on the latent part, whereas the key and value are computed on the entire context. This results in a cross attention in the first layer $\in\mathbb{R}^{l\times d}$ where $l$ is the latent size. The output from the first layer $\in\mathbb{R}^{l\times d}$. The remaining layers operate on the $l\times d$ which is smaller than the $n\times d$ size in a standard Transformer. For Transformers with many layers, the PerceiverAR approach is quite efficient because of the smaller size being operated upon after the first layer. Since the history component of the input is not used in autoregressive modeling, and since it also gets compressed into the latent part after the first layer, we attempt to improve upon these short coming via different architectural enhancements in this paper.
An important advancement in handling very long contexts has been recently proposed in LongLoRA Chen et al. . Even though it computes attention in a sliding window manner, it captures the entire context via a division of the attention heads into two groups. It performs a shift in the attention in the second group to propagate the attention information. We use this intriguing concept in a different way in enhancing the PerceiverAR architecture to accomplish same goals as LongLoRA in terms of attention efficiency.
1. We propose three enhancements to the PerceiverAR architecture to overcome the loss of information that gets compressed into the latent part after the first layer. Each enhancement has efficiency computation overhead tradeoffs.
2. Inspired by the ideas of LongLoRA, we present a simple overlapping PerceiverAR segmented architecture that achieves the computational efficiency of sliding window but with the entire context being available as the computation flows down the layers of the Transformer.
3. In addition to the efficient computation, the pairwise overlapping segment attention extracts more meaningful context resulting in a high performance architecture, which we empirically verify.
In the following section, we first elaborate on the PerceiverAR architecture before presenting enhancements to its design.
3 PerceiverAR Architecture
@@
The fundamental concept behind PerceiverAR is to split the input sequence into two components which we denote as the history and latent sequences. We denote the input as $x$ with corresponding sequence length $n$. After the tokenization and embedding is carried out, the input can be considered as composed of a history component and a latent component as: $x\in\mathbb{R}^{n\times d}=x_{history}||x_{latent}$, where ||||| | indicates the catenation of two components. We denote the history length as $h$ and the latent length as $l$. After the embedding operation, $history\in\mathbb{R}^{h\times d}$ and $latent\in\mathbb{R}^{l\times d}$. The PerceiverAR computes the query only on the latent part in the first layer of Transformer, while the key and values are computed on the entire sequence length of size $n$. Thus the attention computation in the first layer produces an output of dimension $\mathbb{R}^{l\times d}$:
$Q_{latent}=W_{q}x_{latent}\in\mathbb{R}^{r\times d}$ (4)
@@
Since the output from first layer of Transformer $\in\mathbb{R}^{l\times d}$ the remaining layers do a normal attention on inputs of size $l$, without splitting the input into two parts as done in the first layer. For auto-regressive training, an input of $n$ tokens is used to create $(n-1)$ training pairs, such that the expected output of first token is the second token, and the expected token of inputs up to $(m-1)$ tokens is the $m^{th}$ token. The PerceiverAR uses the history part as a fixed input, hence the autoregressive training can only be done on the latent part of input. Thus, to hide the future tokens in the training of the first layer, the upper triangular part of the attention matrix is set to $\infty$- ∞. The remaining layers operate only on the input size of the latent length, so the triangularization of the attention matrix is done on the square matrix corresponding to the latent part.
The attention architecture of the PerceiverAR is shown in Figure 1(a). The additional details of layer normalization, feed forward network and skip connections are omitted. The attention complexity of PerceiverAR in the first layer of Transformer is $O(l\times n)$ while the remaining layers have a complexity of $O(l^{2})$. This provides a significant reduction in computation, especially when $l<n$ and many layers are used in the Transformer. While the PerceiverAR has been able to accomplish impressive results on NLP benchmarks Hawthorne et al. (2022). It has two main drawbacks which we address in this work:
@@
    Lossy History - The history is implicitly compressed into the latent output of first layer and is not explicitly refined as in a normal Transformer via many attention layers.
We improve upon the above drawbacks and present three different enhancements for better utilization of the history component in the PerceiverAR.
4 PerceiverAR Enhancements
Figure 1: Baseline Hawthorne et al. (2022) and Enhanced PerceiverAR Architecture V1.
Figure 2: Other enhanced PerceiverAR Architectures (V2 and V3).
(a) PAR Block for a sequence of half segments
in the first layer of the LLP architecture.
(b) Second layer of LLP architecture.
(c) After enough layers, previous segments
information is implicitly available in PAR
block’s attention.
(d) LongLoRA Inspired PerceiverAR (LLP) Architecture.
Figure 3: In subfigures (a)-(c) the green rectangles indicates the ATTN block, the red rectangles indicates the PerceiverAR attention. The green blocks are not calculated but contain information because of the PAR block in the previous layer. The attention calculation is done only on the two blocks near the diagonal (indicated by red rectangles). Subfigure (d) shows the LLP architecture.
The baseline PerceiverAR uses the history information explicitly only in the first layer by computing the key and values on the entire input sequence, while the query is computed only on the latent part of the input. To overcome the loss of history information in subsequent layers, we propose the first enhancement where each layer computes two attentions and correspondingly two outputs.
4.1 Enhanced PerceiverAR Architecture V1
In this enhancement, each layer performs two attentions. The first attention is the computation as the PerceiverAR baseline as given by Equations 4 through 7. The second attention is based on the history component of the input and also computes the attention in each layer. The second attention generates an additional output as follows:
@@ 4.1 Enhanced PerceiverAR Architecture V1
Thus each layer in the enhanced V1 architecture is identical. The two outputs corresponding to the latent attention and the history attention are catenated to become the single output and the corresponding input for the subsequent layer. Note that no masking is used in the attention on the history part as this part is not used for autoregressive training. Only the latent attention uses masking. This enhanced architecture V1 is depicted in Figure 1(b).
The overhead in this enhancement is the computation of the attention in the history component of the input. If the history length $h$ is larger than the latent length $l$, then this could be significantly more computations as compared to the baseline PerceiverAR where the subsequent layers after the first layer compute attention only on the latent part. To improve this drawback, we propose enhancement V2 as described in the next subsection.
4.2 Enhanced PerceiverAR Architecture V2
To make the history attention computation more efficient, we refine the V1 architecture by dividing the history component into smaller segments. Thus if the segment size $s$ in the history part is smaller than the latent length $l$, i.e., $s<<l$, the overhead in the history computation in each layer is minimal. Note that the segment-wise attention is carried out within the same segment only, so if the segment size is $s$, the complexity of attention in each segment is $O(s^{2})$. The output corresponding to each segment is catenated to become the history component of the output. The architecture of V2 version is shown in Figure 2(a). We present one more enhancement for propagating the history part in a more efficient manner to the layers after the first layer.
4.3 Enhanced PerceiverAR Architecture V3
In this version, the first layer compresses the history part of the input by projecting it to a smaller length along the sequence dimension. This compresses the history information and this compressed history is then used and refined in all remaining layers. The attention and the computation of outputs in enhanced V3 architecture are shown visually in Figure 2(b) and given as follows. First, the the projection of the history to a compressed length $p$ is carried out only in the first layer:
@@ 4.3 Enhanced PerceiverAR Architecture V3
4.4 Long LoRA Inspired PerceiverAR (LLP) Architecture

### v2

# Efficient Context Propagating Perceiver Architectures for Auto-Regressive Language Modeling

@@ Abstract
One of the key challenges in Transformer architectures is the quadratic complexity of the attention mechanism, which limits the efficient processing of long sequences. Many recent research works have attempted to provide a reduction from the $O(n^{2})$ time complexity of attention to semi-linear complexity. However, it remains an unsolved problem in the sense of maintaining high performance when complexity is reduced. One of the important works in this respect is the Perceiver class of architectures that have demonstrated excellent performance, while reducing the computation complexity. In this paper, we use the PerceiverAR as a basis and explore the design space of different trade-offs between preserving context and reducing attention complexity. To this end, we develop four new architectural paradigms, the best performing of which we denote as the Efficient Context propagating Perceiver (ECP). ECP has two major advantages over the PerceiverAR. First, the ECP architecture overcomes the main drawback of PercieverAR by utilizing both the context and the latent sequences in autoregressive training. Second, the ECP architecture operates with the same attention complexity as LongLoRA, making it computationally efficient. More importantly, via pairwise segment attention, it extracts better information resulting in improved language modeling. Empirically, we demonstrate that the ECP architecture significantly outperforms other state-of-the-art Transformer models on Wikitext-103, PG-19 and sCIFAR-10.
Code available at: https://github.com/MetaMain/ECPTransformer
[…]
123
@@ 1 Introduction
The Transformer architecture has revolutionized the field of artificial intelligence, especially in Natural Language Processing (NLP) [24]. The recent success of Large Language Models (LLMs) such as ChatGPT  [1], Gemini [22], Llama [23, 9], etc. with their comprehension and reasoning capabilities, is a testament to the effectiveness of the Transformer architecture. Prior to Transformers, deep Convolutional Neural Networks (CNNs) had demonstrated impressive results in computer vision applications, however, their performance does not show the same effectiveness when applied to NLP. One of the reasons CNNs are typically less effective in NLP is their limited receptive field, due in part to the convolution operation. The Transformer on the other hand, uses the attention mechanism. This operation measures the pairwise similarity between the words or tokens of the entire input sequence in order to comprehend it.
Attention is one of the major computational operations in Transformer architectures with an $O(n^{2})$ complexity. There has been a great deal of research to reduce the complexity of attention while maintaining high performance. Of particular note is the PerceiverAR [12]. The PerceiverAR efficiently handles long contexts by dividing the input sequence in to two components, history (or the context) and latent. However, in this architecture, the history component of the input is not used in autoregressive learning. An additional drawback of the PerceiverAR is the history component gets compressed into the latent part after the first layer. In this work, we design a new architecture called Efficient Context Propagating Perceiver (ECP). The ECP architecture uses local pairwise segment attention making it more computationally efficient than the original PerceiverAR. More importantly, this leads to better learning as we demonstrate in later sections. The Efficient Context Propagating Perceiver is shown in Figure 1(a) and the visual depiction of the associated algorithm is shown in Figures 1(b-d). As can be seen in the figures, the local pairwise segment attention gets propagated to the entire context as more layers are used in the Transformer. We provide more details on this design in later sections. Overall, our work advances the field of autoregressive language modeling and provides the following contributions:
  1. 1.
    Design of Context Efficient Architectures - We design three different architectures to overcome the loss of information inherent in the history or context part of the PerceiverAR, as it is merged into the latent after the first layer. Each architecture has computation efficiency versus performance trade-offs which we analyze.
  2. 2.
    Efficient Context Propagating PerceiverAR (ECP) - We develop a novel segment attention algorithm that is highly efficient and has similar complexity as LongLoRA. Simultaneously, it achieves implicitly full attention while computing attention only on pairs of overlapping segments in a PerceiverAR style.
  3. 3.
    Empirical Evaluation - Experimental results show that our proposed ECP architecture outperforms SOTA models on various datasets. On Wikitext-103, and PG-19 benchmarks, ECP achieves significantly lower perplexity than equivalent size SOTA models.
Figure 1: Subfigure (a) shows the ECP architecture. In subfigures (b)-(d) the green rectangles indicate the propagation of attention, the red rectangles indicate the PerceiverAR pairwise segment attention. The green blocks are not calculated but contain information because of the PAR block in the previous layer. The attention calculation is done only on the two blocks near the diagonal (indicated by red rectangles).
2 Background and Related Work
@@
For a Transfomer, consider the matrix $Q$, $K$ and $V$ containing rows representing the learnt position encoded (PE) embedding of each token in $d$ dimensions (i.e., $1\times d$). The attention $A=\text{softmax}(QK^{T})$ contains the dot product similarity of each input token with every other token in the input sequence. For an input sequence with $n$ tokens, $Q,K\in\mathbb{R}^{n\times d}$ and attention $A\in\mathbb{R}^{n\times n}$.
The Transformer divides the attention calculation into parallel heads so as to refine the learning, and better comprehend the contextual meaning of the input sequence. Each head computes the attention on a portion of the embedding dimension. The output in each head is computed by further multiplying the attention $A$ with $V$. The canonical Transformer’s operation can be summarized by the following equations [24]. First the output of the $i^{th}$ head is:
@@
For language models, if text generation is the goal, the model is trained in an autoregressive manner, where it learns to predict the next token given an input sequence of tokens. In autoregressive generation, the previously predicted token becomes part of the next input sequence. For NLP models, the training process can be made highly effective by masking the attention matrix in a triangular fashion so that future tokens are not visible. This helps in creating more (input, output) training pairs. From an input training sequence of size $n$, n1n-1 training pairs can be created by simply hiding the next token, one at a time. For NLP classification, the masking of the attention is not needed, as the classification decision is made on the entire input sequence.
@@
Since the attention computation in each head measures the pairwise similarity in the input sequence, its time complexity is $O(n^{2})$ if the sequence length is $n$. With larger NLP models being created operating on longer sequence lengths, with multiple heads in each layer and many layers, the computational costs of Transformer training is becoming an important issue. Considerable research effort is being put into making the attention mechanism more efficient since it is the dominant computation in a transformer. Numerous research papers have proposed ideas to reduce the quadratic time complexity of attention to linear or sub quadratic complexity.
Some of the important works in this respect include: TransformerXL [6], Linformer [25], Longformer [3], Reformer [16], Performer [5], Long-Short Attention [26], and Perceiver [12, 15, 14], among others. Recently State Space Models [11, 7, 8] have drawn considerable research attention producing impressive results in some domains. It is yet to be seen if they can be better alternatives for NLP generative models. Thus, we limit our comparisons to transformer-based architectures in this work. Next we elaborate on the PerceiverAR [12] architecture that we use as a basis to further build and arrive at an efficient and better transformer architecture.
3 Preliminaries - PerceiverAR Architecture
@@
The fundamental concept behind PerceiverAR [12] is to split the input sequence into two components which we term as the history and latent sequences. We denote the input as $x$ with corresponding sequence length $n$. After the tokenization and embedding is carried out, the input can be considered as composed of a history component and a latent component as: $x\in\mathbb{R}^{n\times d}=x_{history}||x_{latent}$, where $||$ indicates the catenation of two components. We denote the history length as $h$ and the latent length as $l$. After the embedding operation, $history\in\mathbb{R}^{h\times d}$ and $latent\in\mathbb{R}^{l\times d}$. The PerceiverAR computes the query only on the latent part in the first layer of Transformer, while the key and values are computed on the entire sequence length of size $n$. Thus the attention computation in the first layer produces an output of dimension $\mathbb{R}^{l\times d}$:
$Q_{latent}=W_{q}x_{latent}\in\mathbb{R}^{l\times d}$ (4)
@@
Since the output from first layer of Transformer $\in\mathbb{R}^{l\times d}$ the remaining layers do a normal attention on inputs of size $l$, without splitting the input into two parts as is done in the first layer. For autoregressive training, an input of $n$ tokens is used to create (n1)(n-1) training pairs, such that the expected output of first token is the second token, and the expected token of inputs up to (m1)(m-1) tokens is the $m^{th}$ token. Since PerceiverAR uses the history part as a fixed input, the autoregressive training can only be done on the latent part of input. Thus, to hide the future tokens in the training of the first layer, the upper triangular part of the attention matrix (corresponding to latent) is set to -$\infty$. The remaining layers operate on the size of the latent length, so the triangularization of the attention matrix is done on the entire square matrix.
The attention complexity of PerceiverAR in the first layer of Transformer is $O(l\times n)$ while the remaining layers have a complexity of $O(l^{2})$. This provides a significant reduction in computation, especially when $l$ is much less than $n$ and many layers are used in the Transformer. While the PerceiverAR has been able to accomplish very good results on NLP benchmarks [12], it has two main drawbacks which we significantly improve in this work, finally leading to an impressive new architecture:
@@
    Lossy History - The history is implicitly compressed into the latent output of first layer and is not explicitly refined as in a normal Transformer in following layers.
We improve upon the above drawbacks and present three different enhancements for better utilization of the history component in the PerceiverAR. This further leads to the best architecture developed in this work.
4 New Architecture Paradigms for Latent History Attention and Efficiency
Figure 2: Baseline [12] and Double Attention PerceiverAR architectures.
The baseline PerceiverAR (as shown in Figure 2(a)) uses the history information explicitly only in the first layer by computing the key and values on the entire input sequence, while the query is computed only on the latent part of the input. To overcome the loss of history information in subsequent layers, we propose the first architectural paradigm where each layer computes two attentions and correspondingly two outputs.
4.1 Double Attention PerceiverAR Architecture
In this enhancement (shown in Figure 2(b)), each layer performs two attention operations. The first attention is the same computation as the PerceiverAR baseline as given by Equations 4 through 7. The second attention computation is based on the history component of the input, and is computed in each layer. The second attention generates an additional output as follows:
@@ 4.1 Double Attention PerceiverAR Architecture
Thus each layer in the double attention PerceiverAR architecture is identical. The two outputs corresponding to the latent attention and the history attention are concatenated to become the single output, and therefore the corresponding input for the subsequent layer. Note that no masking is used in the attention on the history part, as this part is not used for autoregressive training. Only the latent attention uses triangular masking.
The overhead in this architecture is the computation of the attention in the history component of the input. If the history length $h$ is larger than the latent length $l$, then this could be significantly more computations as compared to the baseline PerceiverAR where the subsequent layers after the first layer compute attention only on the latent part. To improve this drawback and reduce the computational overhead, we propose a second architectural paradigm as described in the next subsection.
4.2 Compressed Double Attention PerceiverAR Architecture
Figure 3: Compressed Double Attention PerceiverAR,
To improve the efficiency of the history attention computation, we refine the double attention PerceiverAR architecture. We reduce the attention overhead by having the first layer compress the history part of the input by projecting it to a smaller length along the sequence dimension as shown in Figure 3. This compresses the history information and this compressed history is then used and refined in all remaining layers. We detail the mathematics behind this architecture as follows: First, the projection of the history to a compressed length $p$ is carried out only in the first layer:
@@ 4.2 Compressed Double Attention PerceiverAR Architecture
It is important to note that the compressed double attention architecture is one possible way to achieve efficient attention computations. While a variety of other approaches are possible, our next architectural paradigm demonstrates another intuitive way that efficient computation can be achieved.
4.3 $s$-Split Double Attention PerceiverAR
In this architectural paradigm, efficient computation is achieved by dividing the history component into small segments of size $s$. If the history segment size $s$ is smaller than the latent length $l$, i.e., $s<<l$, then the overhead in the history computation in each layer is minimal. Note that the segment-wise attention is carried out within the same segment only, so if the segment size is $s$, the complexity of attention in each segment is $O(s^{2})$. The output corresponding to each segment is catenated to become the history component of the output. A visualization for the $s$-Split double attention PerceiverAR is shown in Figure 4.
Figure 4: $s$-Split Double Attention PerceiverAR.
The above architecture paradigm leads to a highly efficient information propagation and extraction paradigm that we detail in the next section.
5 Efficient Context Propagating Perceiver (ECP) Architecture
In general, the complexity of computing attention is $O(n^{2})$. In Section 4, we introduced several novel paradigms to propagate the history component more effectively in the PerceiverAR architecture, and also make the history attention computationally efficient. First, in the double attention PerceiverAR, the attention is computed for both the history and latent component. The drawback here is the added computational burden of computing double the number of attentions. One way to reduce this complexity is by dividing the input sequence into disjoint segments, and only computing the attention in each segment itself (the $s$-split double attention PerceiverAR). However, the disadvantage of such an approach is that there is some loss of context due to lack of information flow between disjoint segments. Likewise, the compressed double attention PerceiverAR may also lose information as the history component is compressed in the first layer.